Find the APR, or stated rate, in each of the following cases. (Do not round intermediate calculations and enter your answers as a percent rounded to 2 decimal places, e.g., 32.16. Use 365 days in a year.) 


Stated Rate (APR)Number of Times CompoundedEffective Rate (EAR)

%Semiannually13.75%


Monthly9.75


Weekly11.25


Daily9.25
Answer #1

APR = m[(1 + EAR)1/m – 1]

APR = 2[(1.1375)1/2 – 1] = 0.1331 or 13.31%

APR = 12[(1.0975)1/12 – 1] = 0.0934 or 9.34%

APR = 52[(1.1125)1/52 – 1] = 0.1067 or 10.67%

APR = 365[(1.0925)1/365 – 1] = 0.0885 or 8.85%